Jim is 21 years older than Laddie.The sum of their ages is 121.How old is Jim?How old is Laddie
Answer by TimothyLamb(4379)

J = L + 21
J + L = 121
---
J + L = 121
L + 21 + L = 121
2L = 121 - 21
2L = 100
---
L = 50
J = 71
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php